OpenStudy (anonymous):

write this quadratic equation in standard form. then use the quadratic formula to find the solutions. x squared +4x-5=0. please show work:)

jhonyy9 (jhonyy9):

x^2 +4x -5=0 D=16+20=36 x_1,2=(-4+/- sqrt36)/2 = (-4+/- 6)/2 = 2(-2+/- 3)/2 = -2+/-3 = -5 and 1

jhonyy9 (jhonyy9):

x^2 +4x -5=(x+5)(x-1)

jhonyy9 (jhonyy9):

ok ?